OpenStudy (anonymous):

Need help! Will fan and medal!! A circle is centered at (−3, 2) and has a radius of 2. Which of the following is the equation for this circle? (x + 3)2 + (y − 2)2 = 2 (x − 3)2 + (y + 2)2 = 4 (x + 3)2 + (y − 2)2 = 4 (x − 3)2 + (y + 2)2 = 2

OpenStudy (anonymous):

@Ares0706 @Nnesha

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

I need help

OpenStudy (anonymous):

@sourwing @jrx120637 @mathstudent55 @surjithayer @prowrestler

OpenStudy (anonymous):

\[if~ centre~ of~ circle ~is~\left( h,k \right)\] then eq. of circle is \[\left( x-h \right)^2+\left( y-k \right)^2=r^2\] where r is the radius of circle.